Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

Set of numbers

Expert replies
by alex.gellatly » Mon Jul 30, 2012 11:55 pm
This is from 300+ Math questions and best solutions:

6. There are two set each with the number 1, 2, 3, 4, 5, 6. If randomly choose one number from each set, what is the probability that the product of the 2 numbers is divisible by 4?

[spoiler]Solution:
Picking 2 numbers from each set : 6c1*6c1=36
Favorable outcomes = 15
(1,4)
(2,2),(2,4),(2,6)
(3,4)
(4,1),(4,2),(4,3),(4,4),(4,5),(4,6)
(5,4)
(6,2),(6,4),(6,6)
Therefore Required probability = 15/36.[/spoiler]

I was wondering if there was a different way than simply picking favorable outcomes. Also.. I am a little confused why we count both (2,4) and (4,2), but then only (2,2) once. Shouldn't we count (2,2) twice from the same logic we count (2,4) twice?

Thanks
A useful website I found that has every quant OG video explanation:

https://www.beatthegmat.com/useful-websi ... tml#475231
Join the discussion
Source: — Problem Solving |

by armand_h » Tue Jul 31, 2012 12:45 am
P(x4) = 1- P(not x4)
P (not x4) is equal to:
probability that the 1st number is odd and the second is not 4
or
the 1st number is even (but not 4) and the second is odd
=3/6*5/6 + 2/6*3/6
=21/36

P(x4)=1-21/36=15/36
Join the discussion

by eagleeye » Wed Aug 01, 2012 10:18 am
alex.gellatly wrote:This is from 300+ Math questions and best solutions:

6. There are two set each with the number 1, 2, 3, 4, 5, 6. If randomly choose one number from each set, what is the probability that the product of the 2 numbers is divisible by 4?

[spoiler]Solution:
Picking 2 numbers from each set : 6c1*6c1=36
Favorable outcomes = 15
(1,4)
(2,2),(2,4),(2,6)
(3,4)
(4,1),(4,2),(4,3),(4,4),(4,5),(4,6)
(5,4)
(6,2),(6,4),(6,6)
Therefore Required probability = 15/36.[/spoiler]

I was wondering if there was a different way than simply picking favorable outcomes. Also.. I am a little confused why we count both (2,4) and (4,2), but then only (2,2) once. Shouldn't we count (2,2) twice from the same logic we count (2,4) twice?

Thanks
Yes, we can do it without counting the individual cases:

The product will be a multiple of 4 in two cases:

1. Both numbers are even. No. of such cases = 3(choices of even for 1st)*3(choices of even for second) = 3*3 = 9
2. One of the numbers is odd, and other is a 4. Number of such cases = 2(selecting which set is the 4 chosen from)*1 (the choice for 4)*3(the choice for odd numbers) = 2*1*3 = 6

Hence probability = (9+6)/(6*6) = 15/36 = 5/12.

:)
Join the discussion